# Microstructural characterization to reveal evidence of shock deformation in a Campo del Cielo meteorite fragment

### v1

@@ Abstract
The study of meteorites and their microstructures is a topic which spans multiple fields of research, such as earth & planetary sciences, meteoritics, geosciences, and materials science. For materials scientists and engineers, the extreme and unusual conditions which these microstructures form allow for insight into materials which would exist at the edge of our thermomechanical processing abilities. One such microstructure found in low-shock event iron meteorites is Neumann bands. These bands are an array of lenticular deformation twins that form throughout the Fe-Ni matrix with numerous intersections, resulting in many high stress and strain regions within the material’s surface. The existence of these regions and the shocks that formed them encourage atypical strain accommodating mechanisms and structural changes of the material. However, direct investigation of the deformation twin intersections and the microstructural behaviour in and around these regions has been limited. In this work, investigation of these regions in a Campo del Cielo meteorite fragment, with electron backscatter diffraction (EBSD) and forescatter electron (FSE) imaging, revealed two primary findings: high-intensity pattern doubling mirrored across the {110} band at twin-twin intersection and microband formation across the sample surface, which suggest multilayer twinning and constraint of the crystal structure at points of twin-twin intersection. Microbands were found to form along the {110} plane and in regions near Neumann bands. The simultaneous existence of Neumann bands (microtwins) and microbands is presented here for a BCC material, and it is believed the Neumann band and microbands formed during different types and/or shock events from one another. The presence of both Neumann bands and microbands within a BCC iron meteorite is previously unreported and may be valuable in furthering our understanding of shock deformation within iron-based materials.
1 Introduction
@@
Meteorites and their fragments provide insight into the formation of materials on earth and in space. In brief, a meteorite is a meteor that survives its passage through the atmosphere of the earth and strikes the ground. These meteorites often contain very interesting microstructures that are uncommonly found on earth, as they have unique chemistry and thermomechanical processing histories (e.g. very long cooling periods) that depend on the meteor’s formation, movement through space, and location of impact. When the meteor becomes a meteorite, it may experience short-lived rapid shock which can result in some unusual characteristic microstructural features as deformation occurs at very high strain rates (estimated between and ) [1], often with internal aftershocks. Therefore, the ultimate microstructure of a meteorite fragment is dependent on the impact and thermal histories of the host meteor and the resulting meteorite [2, 3], in a similar manner to how the microstructure of an engineering component is influenced by the manufacturing process. These high strain rate microstructures, resulting from impact shocks, can be characterized to understand these otherworldly materials, and provide insight into the extremes of our accessible thermomechanical processing envelope.
@@
The present manuscript focuses on the Campo del Cielo meteorite, which is an iron-rich Fe-Ni metallic meteorite. Major phases in the Fe-Ni system include kamacite () and taenite (). The kamacite phase is closely related to the ferrite phase in steel, with a body centred cubic (BCC) crystal structure and a chemistry that is Fe-rich with typically 3.3-5.8 wt% Ni [4, 5]. The taenite phase is closely related to the austenite phase in steel, with a face centered cubic (FCC) crystal structure and a chemistry that is Fe-rich with typically 31 to 54 wt% Ni [4, 5]. Additionally, the presence of other elements, such as P, S, and Co, results in the precipitation of different minerals within the iron matrix. A common mineral precipitate that is found is schreibersite (also known as rhabdite), an iron-nickel phosphide mineral with a tetragonal crystal structure and appears as a blocky precipitate. This precipitate is found to form in preferential orientations, with the orientation relationship given in Eqn. 1 and 2 [6]. Iron meteorites are also known to have other microstructural features which form in preferential orientations, such as Neumann bands.
@@
Single slip deformation in BCC metals is also known to result in the formation of another visually unique microstructural feature called microbands [14, 15]. Microbands appear as linear bands, coincident with the {110} planes in BCC metals [16], and form initially from strain-induced primary slip planes with independent dislocations. Shear stress acting on these slip planes then results in a polarized arrangement of the dislocations and the formation of double dislocation walls. A more in-depth explanation of the microband formation mechanism is given by Huang and Gray [15]. While these bands are mechanistically independent from Neumann bands, they may share traces of {110} with Neumann band traces of {112} along the [110] orientation [16]. With that said, microbands and Neumann bands have been previously observed within iron and steel materials separately [17, 18] and in lab-based impact tests in stainless steel "crater-related microstructures are dominated by deformation microtwins with a few intermixed microbands" [19]. However, we are not aware of previously documented simultaneous existence of these microstructural features for iron meteorites.
@@ 2.1 Material Preparation
A meteorite fragment from the Campo del Cielo meteorite group (purchased from Rocks Gems Canada, Whistler, BC) was metallographically polished using 400, 600, 800, and 1200 grit silicon carbide for 10 minutes each, with the sample rotated after 5 minutes per grit. This was followed by polishing with a and diamond suspension, and a colloidal silica suspension for 6 minutes each (again rotating by halfway through). Lastly, broad ion beam (Ar-ions) polishing was performed using a Gatan PECS II system for 1 hour at tilt, , 1 rpm, and no modulation.
@@ 2.2 Material Analysis
Electron microscopy was performed on a Tescan AMBER-X. Secondary Electron (SE) images were obtained using a beam current, accelerating voltage, and a working distance of 13-14 . EBSD was performed with an Oxford Instruments Symmetry S2 detector using beam current and accelerating voltage, with the sample loaded into a pre-tilted holder. EBSD acquisition was performed three times with increased magnification on regions of interest (ROI). The first ROI was imaged with an acquisition speed of , a resolution of 156x, and an exposure time of , the second ROI at , 1244x, and and the third ROI at , 622x, and . Further Forescatter electron (FSE) imaging was performed with an increased dwell time of at a working distance of , , and .
@@
The sample was mounted flat for energy dispersive X-ray spectroscopy (EDS/EDX), and area, point and mapping was performed using an Oxford Instruments EDS analysis using an Ultim Max 170 detector. Analysis was performed at , , and a working distance of , using Oxford Instruments Aztec with the Pulse Pile up correction.
@@ 3 Results
Fig. 1 shows an overview of the area studied, and reveals that the meteorite fragment was found to be composed of kamacite with a large number of cuboid schreibersite precipitates, as confirmed with EBSD (crystallography) and EDS (chemistry). EDS results were as expected for kamacite and schreibersite. The kamacite matrix was composed of 88.8-94.0 wt% Fe and 5.8-6.0 wt% Ni, any remainder was attributed to be carbon. The schreibersite precipitates had elemental composition of 42.6-45.2 wt% Fe, 35.0-36.5 wt% Ni, 13.6-12.3 wt% P, and 8.8-5.9 wt% C. The map sum elemental composition of the meteorite fragment was 88.2 wt% Fe, 5.9 wt% Ni, 5.8 wt% C, 0.1 wt% P. The carbon map is not provided in Fig. 1 as it was found dispersed across the sample and may be from additional surface contamination.
Neumann bands were seen to be interlaced over the entire surface with crystal orientations being consistent amongst parallel bands. The boundaries of these bands were found to be <111> twins with a misorientation. Areas of elevated lattice rotation gradient (i.e. strain gradient) can be seen around the intersection point of the Neumann bands, shown by the slight changes of orientation in the IPF plots. These regions are most visible in the IPF-X of Fig. 1 and further investigation was performed at the isolated twin tip region in Fig. 2.
@@
A further region of interest was identified near the isolated twin tip, shown in Fig. 3. The region in question was identified to have a high KAM value and a large number of dislocations within a small twin intersection. Initial indexing of this location failed and direct investigation resulted in the discovery of the Electron Backscatter Patterns (EBSPs) seen in Fig. 3.
SE imaging also lead to the discovery of a seemingly ‘rippled’ surface as seen in Fig. 4. These ‘ripples’ were seen to run parallel to one another, with a consistent spacing of approximately . The ripples were found to be present over large distances and were shifted when near Neumann bands.
@@ 4 Discussion
This Campo del Cielo fragment was found to be covered in Neumann bands, with numerous intersections points between these bands. Closer investigation of isolated Neumann bands/twins revealed a network of dislocations (Fig. 2) from an isolated twin tip to the Neumann band twin present above it and a higher WBV magnitude, and therefore dislocation density, along the boundary of the lower Neumann band.
The dislocation density on the lower Neumann band varied along the boundary, with areas of high density dotted along the upper side of the twin grain boundary. It is noted that the increased density is not seen in the IPF-Y and Z (Fig. 2) on the lower side of the grain boundary. This may be a result of the grain growth mechanism for BCC deformation twins. As found by Zhong et al. for BCC tantalum, the twin growth in the BCC metals may occur through a nucleation-controlled mechanism or a growth-controlled mechanism [11]. The later mechanism occurs with the twin grain having a coherent twin boundary (CTB) on one side and an incoherent twin boundary on the other. The widening of the twin occurs on the CTB side and results in a gradual build up of lattice strains and strain energies on the incoherent side until the energy required for nucleation of another CTB to occur is overcome.
This growth-controlled mechanism is believed to be present in the lower Neumann band, with the CTB being the lower side of the band and the incoherent boundary being the upper side. High strain energies and raised dislocation density are seen on the upper side and indicates that the twin grain has not reached the critical strain energy required to initiate further growth. From Fig. 2 it can also be seen that these high strain regions are present at locations when the band is thinner, but adjacent to thicker regions. While the growth occurs rapidly once the energy barrier is overcome [11], it may be that when this barrier is reached that it is not done uniformly across the twin grain boundary.
@@
The isolated twin tip with the network of dislocations does not exhibit this same behaviour. Rather, the twin tip does not have a CTB and incoherent boundary but instead exhibited a lower KAM and WBV magnitude along its body, with increased magnitudes at and near the tip. It is possible that the increased dislocation density near the tip is a precursor to new grain boundary formation through a nucleation-controlled mechanism. Similar behaviour has been simulated previously by Sainath and Nagesha for the nucleation and initiation of a BCC iron twin [27].
However, if this twin tip were to further propagate it would be outward from the tip point, rather than upward toward the Neumann band in the upper right of the WBV magnitude in Fig. 2.
@@
Further analysis of the smallest region of interest lead to the discovery of the unusual EBSP shown in Fig. 3 in this region of very large deformation near the edge of the twin-twin interface. The twin and matrix EBSPs show that there is a common {110} band within these diffraction patterns, and the unindexed region (in region 3 of the map) has a pattern which contains a single diffraction band superimposed upon a noisy ’background’ (which has some evidence of other diffracting features). In general, in the presence of a strain gradient within the interaction volume, it is well known and well understood that the EBSD-pattern can blur and become harder to index. In this unusual example, all but one band has blurred, and this indicates that the strain gradient for this volume is highly restricted and there is crystalline material with a (relatively) undisturbed {110} plane and one could imagine that this is the extreme case of ’pattern doubling’ that could be related to either a very high strain gradient or multilayer twinning [28] in this localized region. This motivated the HR-EBSD based GND analysis, which supports the idea that there is a localized strain gradient in this location due to the presence of a very high GND density near this region.
@@
A multilayer twin is a section of the material where multiple twin deformations are stacked atop one another. These multilayer twins are denoted by the number of packing layers between the twins, with density of twin increasing as the number is lower [28]. For example, a 9-layer twin has 9 layers (ABCABCABC) between it and the twin deformation plane (illustration provided in [28]). When the electron beam direction is parallel to the twin plane of this stack it results in overlap of EBSP of each twin, causing constructive interference near the twin plane band and increasingly destructive interference further away [28]. This behaviour has been simulated previously by Li et al. for FCC iron down to 3-layer twins. The simulated EBSP presented by Li et al. appears similar in nature to the unusual EBSP seen in this work, however the EBSP seen here experiences greater amounts of interference (constructive and destructive). Therefore it is hypothesized that if multilayer twinning is present in this EBSP, then it is at a greater degree than 3-layer.
@@
Constraint of the crystal structure may also play a part in the high {110} presence seen in the EBSP. Intersection of the two twin bands results in the generation of shear stress and movement through specific slip systems. Those being {110}, {112}, and {123} for BCC metals, in order of increasing rarity [29]. The most common slip plane, {110}, is known to be related to the formation of Neumann bands, with {110} acting as a twin plane for the formation of Neumann bands [7]. It is suggested that the twin-twin intersection and the shear stress involved results in slip along the {110} plane, constraining the microstructure and its dislocations. Thereby, causing a high intensity {110} band to be seen at the site of exit for the twin-twin intersection.
@@
SE imaging of this region also showed an interesting surface phenomena. A ripple-like appearance was seen (Fig. 5) that seemed to travel along planes similar to the large Neumann bands. Surface traces on the {110} and {112} pole figures were performed, given the known relation of these planes to Neumann bands [8, 7]. This revealed an overlap with the {110} plane and ‘wavefront’ of the ripple, and for the {112} plane with the ripple-ripple interface. Additional ripples were seen over large regions of the sample, shown in Fig. 4, that have the same ‘wavefront’, but are not seen to have the same ripple-ripple interfaces as in Fig. 5.

### v2

@@ Abstract
For materials scientists and engineers, the extreme and unusual conditions which meteorites and their microstructures form allow for insight into materials which would exist at the edge of our thermomechanical processing abilities. One such microstructure found in low-shock event iron meteorites is Neumann bands. These bands are arrays of lenticular deformation twins that form throughout the Fe-Ni matrix with numerous intersections, resulting in many high stress and strain regions within the material’s surface. These regions and the shocks that formed them encourage atypical strain accommodating mechanisms and structural changes of the material. However, investigation of the deformation twin intersections and the microstructural behaviour in and around these regions has been limited. In this work, investigation of these regions in a Campo del Cielo meteorite fragment, with electron backscatter diffraction (EBSD) and forescatter electron (FSE) imaging, revealed two primary findings: high-intensity pattern doubling mirrored across the {110} band at twin-twin intersection and microband formation across the sample surface, suggesting multilayer twinning and constraint of the crystal structure at twin-twin intersection points. Microbands were found to form along the {110} plane and in regions near Neumann bands. The simultaneous existence of Neumann bands (microtwins) and microbands is presented here for a BCC material, and it is believed the Neumann band and microbands formed during different types and/or shock events from one another. The presence of both Neumann bands and microbands within a BCC iron meteorite is previously unreported and may be valuable in furthering our understanding of shock deformation within iron-based materials.
keywords:
Iron-meteorite, Iron BCC, EBSD, Microbands, Neumann bands, Pattern doubling1 Introduction
@@
Meteorites and their fragments offer valuable insights into the processes governing material formation both on Earth and in extraterrestrial environments. In brief, a meteorite is a meteor that survives its passage through the atmosphere of the earth and strikes the ground. These meteorites often contain very interesting microstructures that are uncommonly found on earth, as they have unique chemistry and thermomechanical processing histories (e.g. very long cooling periods) that depend on the meteor’s formation, movement through space, and location of impact. When the meteor becomes a meteorite, it may experience short-lived rapid shock which can result in some unusual characteristic microstructural features as deformation occurs at very high strain rates (estimated between and ) [1], often with internal aftershocks. Therefore, the ultimate microstructure of a meteorite fragment is dependent on the impact and thermal histories of the host meteor and the resulting meteorite [2, 3], in a similar manner to how the microstructure of an engineering component is influenced by the manufacturing process. These high strain rate microstructures, resulting from impact shocks, can be characterized to understand these otherworldly materials, and provide insight into the extremes of our accessible thermomechanical processing envelope.
@@
The present manuscript focuses on the Campo del Cielo meteorite, which is an iron-rich Fe-Ni metallic meteorite. Major phases in the Fe-Ni system include kamacite () and taenite (). The kamacite phase is closely related to the ferrite phase in steel, with a body centred cubic (BCC) crystal structure and a chemistry that is Fe-rich with typically 3.3- [4, 5]. The taenite phase is closely related to the austenite phase in steel, with a face centered cubic (FCC) crystal structure and a chemistry that is Fe-rich with typically 31 to [4, 5]. Additionally, the presence of other elements, such as P, S, and Co, results in the precipitation of different minerals within the iron matrix. A common mineral precipitate that is found is schreibersite (also known as rhabdite), an iron-nickel phosphide mineral with a tetragonal crystal structure and appears as a blocky precipitate. This precipitate is found to form in preferential orientations, with the orientation relationship given in Eqn. 1 and 2 [6]. Iron meteorites are also known to have other microstructural features which form in preferential orientations, such as Neumann bands.
@@
Single slip deformation in BCC metals is also known to result in the formation of another visually unique microstructural feature called microbands [14, 15, 16]. Microbands appear as linear bands, coincident with the {110} planes in BCC metals [17], and form initially from strain-induced primary slip planes with independent dislocations. Shear stress acting on these slip planes then results in a polarized arrangement of the dislocations and the formation of double dislocation walls. A more in-depth explanation of the microband formation mechanism is given by Huang and Gray [15]. While these bands are mechanistically independent from Neumann bands, they may share traces of {110} with Neumann band traces of {112} along the [110] orientation [17]. With that said, microbands and Neumann bands have been previously observed within iron and steel materials separately [18, 19] and in lab-based impact tests in stainless steel "crater-related microstructures are dominated by deformation microtwins with a few intermixed microbands" [20]. However, we are not aware of previously documented simultaneous existence of these microstructural features for iron meteorites.
@@ 2.1 Material Preparation
A meteorite fragment from the Campo del Cielo meteorite group (sourced from Rocks Gems Canada, Whistler, BC) was metallographically polished using 400, 600, 800, and 1200 grit SiC for 10 minutes each, with the sample rotated after 5 minutes per grit. This was followed by polishing with a and diamond suspension, and a colloidal silica suspension for each (again rotating by halfway through). Lastly, broad ion beam (Ar-ions) polishing was performed using a Gatan PECS II system for 1 hour at tilt, , , and no modulation.
@@ 2.2 Material Analysis
Electron microscopy was performed on a Tescan AMBER-X plasma focused ion beam-scanning electron microscope (pFIB-SEM). Secondary Electron (SE) images were obtained using a beam current, accelerating voltage, and a working distance of 13-14 . EBSD was performed with an Oxford Instruments Symmetry S2 detector using beam current and accelerating voltage, with the sample loaded into a pre-tilted holder. EBSD acquisition was performed three times with increased magnification on regions of interest (ROI). The first ROI was imaged with an acquisition speed of , a resolution of 156x, a step size of and an exposure time of , the second ROI at , 1244x, and , and the third ROI at , 622x, and . Further Forescatter electron (FSE) imaging was performed with an increased dwell time of at a working distance of , , and .
@@
The sample was mounted flat for energy dispersive X-ray spectroscopy (EDS/EDX), and area, point mapping was performed using an Oxford Instruments EDS analysis using an Ultim Max 170 detector. Analysis was performed at , , and a working distance of , using Oxford Instruments Aztec with the Pulse Pile up correction.
Simulated EBSP were created through reprojection of dynamical pattern simulations. DynamicS was used to generate a reference many beam dynamical EBSD pattern simulation, using a BCC-iron crystal structure (a = ). These simulations were based upon the work of Winkelmann et al. [27] and included reflectors with a lower limit of dhkl = , or the intensity of the bands being of the strongest reflector. The reference sphere was imported into AstroEBSD [28, 29] within MATLAB and the experimental pattern was reprojected using the same pattern centre and crystal orientation as in Fig. 3 c).
@@ 3 Results
Fig. 1 shows an overview of the area studied, and reveals that the meteorite fragment was found to be composed of kamacite with a large number of cuboid schreibersite precipitates, as confirmed with EBSD (crystallography) and EDS (chemistry). The schreibersite precipitates were found to have an average size of , with precipitate sizes ranging from 15 to . EDS results were as expected for kamacite and schreibersite. The kamacite matrix was composed of 88.8- and 5.8-, any remainder was attributed to be carbon. The schreibersite precipitates had elemental composition of 42.6-, 35.0-, 13.6-, and 8.8-. The map sum elemental composition of the meteorite fragment was , , , . The carbon map is not provided in Fig. 1 as it was found dispersed across the sample and may be from additional surface contamination.
Neumann bands were seen to be interlaced over the entire surface with crystal orientations being consistent amongst parallel bands. The boundaries of these bands were found to be <111> twins with an average misorientation. Areas of elevated lattice rotation gradient (i.e. strain gradient) can be seen around the intersection point of the Neumann bands, shown by the slight changes of orientation in the IPF plots. These regions are most visible in the IPF-X of Fig. 1 and further investigation was performed at the isolated twin tip region in Fig. 2.
@@
Furthermore, the component maps of the WBV (i.e. WBV-X, WBV-Y, WBV-Z) show the polarization of the dislocation field associated with each twin variant. For example, the large horizontal twins (twin 2) shows that the WBV is dominated by a vector that points in Y and Z (i.e. has limited X component). The two fields that extend away from this twin, one horizontal and associated with the lower twin-matrix interface and the second running diagonally, are of opposite sign, which is important to ensure that there is closure of the deformation field further away from the twin termination site. The other twin variants show a similar effect that starts at the termination of the twin and extend into the matrix, but the WBV vector is, of course, different (as related to the twin-matrix misorientation for each variant).
A further region of interest was identified near the isolated twin tip, shown in Fig. 3. The region in question was identified to have a high KAM value and a large number of dislocations within a small twin intersection. Initial indexing of this location failed and direct investigation resulted in the analysis of Electron Backscatter Patterns (EBSPs) at the investigated location seen in Fig. 3.
Sum patterns of the BCC-iron EBSP were simulated to evaluate the potential of multilayer twinning at the twin-twin interaction point. The individual common/invariant plane {110} normal vector was calculated using the indexed patterns (obtained from pattern matching with the matrix pattern). A two pattern sum was generated through rotation of the matrix orientation by about this normal vector, resulting in a second pattern that was similar in orientation to the twin orientation but with an exact overlap of the invariant plane (as opposed to a propagation of uncertainty from using pattern matching of both the parent and twin experimental patterns). The original simulation and this twin simulation were summed together and the intensities normalized, resulting in the pattern given in Fig. 3 i). Simulation and evaluation of possible constrained lattice rotation was performed in a similar manner. For the constrained lattice rotation, a simulation of the reference pattern was rotated about the same normal vector 360 times in increments, and the sum pattern of this rotation series was calculated and the intensities normalized. The resulting pattern is provided in Fig. 3 h).
FSE imaging also led to the discovery of a seemingly ‘rippled’ surface as seen in Fig. 4. While these ‘ripples’ appear to be deformations on the material surface, they are instead surface contrast resulting from electron channelling effects through the use of FSE imaging. Further discussion of this channelling behaviour is provided in [31, 32, 33]. These ‘ripples’ were seen to run parallel to one another, with a consistent spacing of approximately . The ripples were found to be present over large distances and were shifted when near Neumann bands.
@@ 4 Discussion
This Campo del Cielo fragment was found to be covered in Neumann bands, with numerous intersections points between these bands. The measured thickness of the Neumann bands varied between and , with an average misorientation angle of . Closer investigation of isolated Neumann bands/twins revealed a network of dislocations (Fig. 2) from an isolated twin tip to the Neumann band twin present above it and a higher WBV magnitude, and therefore dislocation density, along the boundary of the lower Neumann band.
The dislocation density on the lower Neumann band varied along the boundary, with areas of high density dotted along the upper side of the twin grain boundary. It is noted that this increased density is not observed in the IPF-Y and Z (Fig. 2) on the lower side of the grain boundary.
This notched appearance on the upper twin boundary has previously been observed by Mahajan and Bartlett [34] for shock deformation twins in molybdenum; however, the cause was undetermined. A mechanism for nanoscale BCC tantalum has been proposed by Zhong et al. which explained a similar difference in the twin boundaries through competing nucleation-controlled and growth-controlled mechanisms [11], but these mechanisms have not been validated for the micrometre and millimetre scales of the Neumann bands.
Rather, the Neumann bands likely initially formed through the slip dislocation interaction mechanism [35, 36], as observed in Fe- and Fe- alloys previously [35]. It is noted that no intersecting slip bands were found during microstructural analysis of the Campo del Cielo fragment. This further corroborates that the presence of slip bands is not required for deformation twin formation [34, 37].
The notched appearance of the band may be a result of discrete stepwise growth of the twin boundary along {112} in a similar fashion to the mechanism described by Jiang et al. [38]. Alternatively, the notched appearance of the twin boundary may be due to the concurrent formation of many smaller deformation twins which subsequently grew together, as seen in high strain rate deformation of bulk-polycrystalline BCC tantalum [39].
@@
The isolated twin tip with the network of dislocations is not seen to have this notched appearance. Rather, the twin tip has a relatively flat boundary with a lower KAM and WBV magnitude along its body, and increased magnitudes at and near the tip. It is possible that the increased dislocation density near the tip is a precursor to new grain boundary formation through a nucleation mechanism. Similar behaviour has been simulated previously by Sainath and Nagesha for the nucleation and initiation of a BCC iron twin [40]. However, if this twin tip were to further propagate it would be outward from the tip point, rather than upward toward the Neumann band in the upper right of the WBV magnitude in Fig. 2.
Comparison of the KAM map and WBV maps in Fig. 2 are interesting as:
The WBV provides higher contrast to show areas of heterogenous deformation, as comparing f) and g).
The decomposed vector components (WBV-X, Y and Z – shown in h-j) reveal that dislocation density is polarized around the intersections and tips of the Neumann bands, including those bands that are not revealed as distinctly indexed features.
The polarization of these fields for each of the similar Neumann band variants is similar, which is consistent with what would be expected for conventional deformation twinning that occurs in more conventional deformation experiments (e.g. as reported in Guo et al. [23]).
In this work, the WBV is used to quickly analyse and describe components of the deformation gradient that can be accessed quickly from the measured maps. Overall, we have presented Fig. 2 as an opportunity to provide new insight into the deformed meteorite sample, as well as providing a comparison of the different modes of analysis that are now routinely available within modern EBSD packages.
@@
Further analysis of the smallest region of interest lead to the discovery of the unusual EBSP shown in Fig. 3 in this region of very large deformation near the edge of the twin-twin interface. The twin and matrix EBSPs show that there is a common {110} band within these diffraction patterns, and the unindexed region (in region 3 of the map) has a pattern which contains a single diffraction band superimposed upon a noisy ’background’ (which has some evidence of other diffracting features). In general, in the presence of a strain gradient within the interaction volume, it is well known and well understood that the EBSD-pattern can blur and become harder to index. In this unusual example, all but one band has blurred, and this indicates that the strain gradient for this volume is highly restricted and there is crystalline material with a (relatively) undisturbed {110} plane and one could imagine that this is the extreme case of ’pattern doubling’ that could be related to either a very high strain gradient or multilayer twinning [41] in this localized region. Further work would be required to qualify the interaction volume size and the minimum number/amount of pattern blurring that is required to try to estimate the strain gradient required for the exact amount of blurring seen in the experiment, which is out of the scope of the present work. To assist in understanding the rest of the deformation in this area, this motivated the HR-EBSD based GND analysis, which supports the idea that there is a localized strain gradient in this location due to the presence of a very high GND density near this region.
@@
A multilayer twin is a section of the material where multiple twin deformations are stacked atop one another. These multilayer twins are denoted by the number of packing layers between the twins, with density of twin increasing as the number is lower [41]. For example, a 9-layer twin has 9 layers (ABCABCABC) between it and the twin deformation plane (illustration provided in [41]). When the electron beam direction is parallel to the twin plane of this stack it results in overlap of EBSP of each twin, causing constructive interference near the twin plane band and increasingly destructive interference further away [41]. This behaviour has been simulated previously by Li et al. for FCC iron down to 3-layer twins. However, the simulated EBSP given in Fig. 3 i) was not seen to match the experimental EBSP (3 g). Rather, the rotation series pattern (3 h) looks more similar to the experimental pattern. Notably, this contains the presence of a bright {110} band with no features of zone axes along the band, as well as blurring of the other bands in the pattern. For the ’double’ pattern shown in Fig. 3 i), the nature of this orientation relationship would leave the {110} band in the top right of the pattern also reproduced exactly in the experiment, which was not observed.
@@
From this simulation, it is believed that constraint of the crystal structure and localized crystal rotation play a part in the high {110} presence seen in the EBSP. Intersection of the two twin bands results in the generation of shear stress and movement through specific slip systems. Those being {110}, {112}, and {123} for BCC metals, in order of increasing rarity [42]. The most common slip plane, {110}, is known to be related to the formation of Neumann bands, with {110} acting as a twin plane for the formation of Neumann bands [7]. It is suggested that the twin-twin intersection and the shear stress involved results in slip along the {110} plane, constraining the microstructure and its dislocations. Thereby, causing a high intensity {110} band to be seen at the site of exit for the twin-twin intersection.
@@
FSE imaging of this region also showed an interesting surface phenomena. A ripple-like appearance was seen (Fig. 5) that seemed to travel along planes similar to the large Neumann bands. Surface traces on the {110} and {112} pole figures were performed, given the known relation of these planes to Neumann bands [8, 7]. This revealed an overlap with the {110} plane and ‘wavefront’ of the ripple, and for the {112} plane with the ripple-ripple interface. Additional ripples were seen over large regions of the sample, shown in Fig. 4, that have the same ‘wavefront’, but are not seen to have the same ripple-ripple interfaces as in Fig. 5.